How would I use the method of substitution for this equation?
-x-y=-2
2x-y=1

Answer by ewatrrr(24785)

-x-y=-2 0r y =+2-x
2x-y=1
2x - (2-x) = 1
3x = 3
x = 1 and y = 1
...
Ordered Pair(1,1) the Solution for this System